# A Perspective on AI-Guided Molecular Simulations in VR: Exploring Strategies for Imitation Learning in Hyperdimensional Molecular Systems

### v1

@@ Abstract
Molecular dynamics (MD) simulations are a crucial computational tool for researchers to understand and engineer molecular structure and function in areas such as drug discovery, protein engineering, and material design. Despite their utility, MD simulations are expensive, owing to the high dimensionality of molecular systems. Interactive molecular dynamics in virtual reality (iMD-VR) has recently been developed as a ‘human-in-the-loop’ strategy, which leverages high-performance computing to accelerate the researcher’s ability to solve the hyperdimensional sampling problem. By providing an immersive 3D environment that enables visualization and manipulation of real-time molecular motion, iMD-VR enables researchers and students to efficiently and intuitively explore and navigate these complex, high-dimensional systems. iMD-VR platforms offer a unique opportunity to quickly generate rich datasets that capture human experts’ spatial insight regarding molecular structure and function. This paper explores the possibility of employing user-generated iMD-VR datasets to train AI agents via imitation learning (IL). IL is an important technique in robotics that enables agents to mimic complex behaviors from expert demonstrations, thus circumventing the need for explicit programming or intricate reward design. We review the utilization of IL for manipulation task domains in robotics and discuss how iMD-VR recordings could be used as datasets to train IL models for interacting with MD simulations and solving specific molecular ‘tasks’. We then investigate how such approaches could be applied to the data structures captured from iMD-VR recordings. Finally, we outline the future research directions and potential challenges of using AI agents to augment human expertise to efficiently navigate vast conformational spaces, highlighting how this approach could provide valuable insight across domains such as materials science, protein engineering, and computer-aided drug design.
(Accepted for presentation at the First Workshop on "eXtended Reality & Intelligent Agents" (XRIA24) @ ECAI24, Santiago De Compostela (Spain), 20 October 2024)
[…]
@@ 1 Introduction
Molecular dynamics (MD) simulations are a powerful tool for studying the structure, dynamics, and interactions of molecular systems. However, generating conformational ensembles and sampling rare events, e.g., protein-ligand binding, remains challenging due to high computational costs and the complexity of the associated energy landscapes[38]. Interactive molecular dynamics in virtual reality (iMD-VR) has recently emerged as a promising approach to address these challenges by leveraging human intuition during real-time MD simulations within an immersive 3D environment[47]. In iMD-VR, users can directly manipulate and steer molecular systems using natural hand motions, applying forces to drive physically-relevant rare events such as conformational changes and ligand binding/unbinding [27]. This human-in-the-loop approach leverages the human’s innate ability for 3D spatial reasoning and manipulation, enabling a user to intuitively explore complex molecular landscapes. Recent studies have demonstrated the efficacy of iMD-VR in recreating crystallographic binding poses for protein-ligand systems[8] [7] and generating important reactive pathways[44]. These interactive simulations capture valuable conformational data that can be challenging to obtain through conventional MD alone, thus offer new opportunities for applications such as training machine learning and investigating reaction mechanisms.
@@
Imitation learning (IL), or learning from demonstration, is a powerful paradigm in artificial intelligence, enabling machines to acquire new skills by observing and mimicking expert behavior[52]. This approach has been particularly influential in the field of robotics, where it has been used to teach robots complex tasks without the need for explicit programming. By observing human demonstrations, robots can learn to perform a variety of actions, ranging from simple manipulations to complex, multi-step procedures[39][32][13][50]. Learning from observation differs from other types of machine learning such as reinforcement learning, where an explicit reward function needs to be defined in advance or fine-tuned during training. IL offers the ability to learn a mapping of observations to actions done by an expert in demonstrations, making it particularly well-suited to domains for which specifying a reward function is challenging or where human expertise can be leveraged. The versatility of IL has also sparked interest in its application beyond robotics, such as in molecular dynamics, where it could potentially streamline the process of simulating and understanding complex molecular interactions.
@@
IL often requires a large number of demonstrations to effectively learn a policy, especially for complex tasks. Collecting a sufficiently large and diverse dataset of human demonstrations can be challenging for various reasons [50]. One way to think of it is that human behavior is often multi-modal—there are many valid ways to perform a task. Standard imitation learning approaches may average out these modes and learn a sub-optimal policy. Capturing and replicating diverse human behaviors is an open challenge [15], mainly, because robustly capturing all relevant aspects of human demonstrations can be difficult due to sensor limitations, occlusions, etc [21]. This is especially true when collecting data outside of lab settings ‘in the wild’. Therefore there is a growing need for more large-scale, open datasets of human demonstrations on standardized tasks in order to facilitate reproducible research and benchmark imitation learning algorithms[15][48].
@@
Virtual reality (VR) presents a novel and immersive platform for enhancing the capabilities of IL, particularly within the field of iMD-VR where VR is combined with high-performance computing to provide an interactive environment in which researchers can manipulate molecular structures in real-time. The intuitive and engaging nature of VR could revolutionize the way scientists interact with molecular simulations, making it easier to collect data, hypothesize, and test the dynamics of molecular systems[8][42].
@@
This paper aims to explore IL’s current applications in various domains, including key concepts like behavioral cloning and generative adversarial imitation learning. Particular focus is given to its potential in molecular dynamics using VR. We aim to provide a comprehensive review of the existing literature on IL, identify the benefits and challenges associated with its use, and propose innovative ways in which VR could serve as a platform for data creation and collection in MD. By bridging the gap between IL and VR, we hope to open new avenues for research and application in the field of MD, ultimately contributing to advancements in scientific understanding and technological development.
@@ 2.1 Molecular vizualisation
Virtual reality (VR) is revolutionizing the way researchers interact with and visualize molecular structures. VR provides researchers with natural, intuitive 3D interfaces to view and interact with complex molecular structures in way that is not facilitated with traditional 2D interfaces. This can enhance the researcher’s understanding of complex 3D molecular arrangements and interactions, which is essential for enabling research insight. Furthermore, VR can enhance scientific collaboration by providing shared virtual environments, which may even be accessible over the internet, thus enabling collaboration across physical distances.
There exist several programs for the visualization of molecular simulations in VR, e.g., UnityMol[10], and the commercial software Nanome[1]. Nanome provides a collaborative virtual environment in which users can visualize and manipulate molecular structures in stereoscopic 3D. Researchers can analyze the spatial arrangement of molecules, measure distances between atoms, and dock ligands into protein binding pockets using natural hand gestures [18]. Other examples of software include ProteinVR [2] and Molecular Rift [26]. ProteinVR is a web-based application that works across desktop, mobile, and VR platforms, democratizing access to structural biology in 3D. Molecular Rift provides controller-free manipulation of molecules using intuitive hand gestures.
By immersing users in 3D virtual environments, VR enables intuitive exploration of complex biomolecular systems that traditional 2D screens do not facilitate. The application of VR to molecular simulations unlocks several key benefits. First, it provides researchers with natural, intuitive 3D interfaces to view and interact with complex molecular structures. Second, by coupling interactive molecular dynamics with VR, scientists can manipulate molecular systems and observe the effects in real-time, potentially uncovering new mechanistic insights. Third, VR enables collaborative drug design and molecular modeling in shared virtual environments, enhancing scientific teamwork. Finally, the stereoscopic depth perception and wide field of view in VR leads to an enhanced spatial understanding of 3D molecular arrangements.[5]
@@ 2.2 Interactive molecular simulations
Recent advancement in computational power and improving performance of graphical processing units has facilitated not only the visualization of molecules in VR, but also real-time interactivity. One prominent example is Narupa, an open-source program developed by Glowacki et al.[14] for performing interactive molecular dynamics in virtual reality (iMD-VR). Using VR controllers, Narupa users can apply forces directly to MD simulations in real-time to drive important chemical events such as ligand binding and conformational changes.
@@
Figure 1: Showing an iMD-VR user docking and undocking the drug oseltamivir from the H7N9 neuraminidase protein [27]
@@
An example of this is demonstrated in Figure 1. The top panel illustrates a researcher using iMD-VR to explore the binding and unbinding pathways of the drug oseltamivir (shown in magenta) with the H7N9 neuraminidase protein. The bottom panel shows snapshots of the molecular system at three timepoints in the simulation: (1) oseltamivir bound to the active site of neuraminidase, (2) the molecular system after the researcher has undocked oseltamivir from the binding pocket of neuraminidase, and (3) the protein-ligand complex, where oseltamivir has been re-docked by the researcher after interactively exploring potential binding modes.
2.3 Data structure in NanoVer
@@
In essence, a molecular simulation is a time series consisting of a set of frames of the atomic positions of a molecular system, which can be viewed frame-by-frame as a ‘trajectory’. In iMD-VR, these trajectories are generated and visualised on-the-fly. Each frame contains information about the system, e.g. the temperature and energy, and the atoms contained within it, e.g. their position and element type. NanoVer (previously known as Narupa) uses a key-value system to store and communicate these frames.
NanoVer streams data in two dictionaries: (a) the frames, containing information about the simulation, such as the atomic positions; and (b) the ‘shared state’, consisting of synchronised information about, e.g., avatar positions and user interactions with the simulation. The NanoVer server can record these data, thus enabling post-hoc analysis and playback of sessions. These recordings can be loaded onto the server, which then sends the recorded streams (synchronised using timestamps) to the clients as if they were real-time simulation streams. In this case, NanoVer acts purely as a molecular visualiser, affording the user control over the position/rotation/scale of the simulation and providing typical playback features (play/pause/restart), though naturally the user cannot apply forces to the molecular system. NanoVer recordings can be imported and analysed using a Python script with the MDanalysis module[33][25].
2.4 Types of molecular simulations
NanoVer provides several molecular simulations that any user can load, or users may import their own OpenMM systems. One of the prototypical examples that we use for NanoVer demonstrations is the simulation of a methane molecule and a carbon nanotube, a molecular system relevant to the study of biomolecular channels that act as molecule-selective filters[17]. In this simulation, players can simulate the action of the nanotube as a biomolecular channel by threading the methane through the nanotube (Figure 2). The molecular system comprises 65 atoms: 60 carbons for the nanotube (labelled C1–C60), and 1 carbon and 4 hydrogens for the methane (labelled C61 and H1–H4). Table 1 shows some example data collected for this task. The output trajectory file ’.traj’ is a binary file containing all data from the molecular simulation. This was converted into a .csv file and the relevant data was filtered for simpler access and processing. The resulting dataframe had 4 columns: atom name, time, coordinates, and user forces. Here ‘name’ is the atom’s label, ‘time’ is the frame index, ‘coordinates’ contains the (x,y,z) positions in nanometers, and ‘forces’ contains the (x,y,z) components of the forces (Fx, Fy, Fz) applied by the user on the specified atom.
Figure 2: Threading Methane through the Nanotube)
@@ 2.4 Types of molecular simulations
Table 1: Dataframe example of the first frame from a recording for the nanotube task
atom name time coordinates user forces
@@
For analysis purposes, the trajectory of the C61 atom from the methane molecule can be plotted as seen in Figure  3. Each subplot represents a distinct attempt to thread the methane molecule through the nanotube.
Figure 3: Atom C61’s trajectory for Nanotube task
A more sophisticated task is the one shown in Figure 4. This task involves tying a knot using a 17 alanine molecule. The user is presented with the molecule in an untied form and asked to form a knot using the movement of both ends of it. Both given examples are demonstration systems involving only small molecules. The software can be used on larger, more complex molecular systems such as protein systems.
@@ 3.1 Recent works in literature
In recent years, imitation learning in agents and multiagent systems has seen significant advancements. One notable contribution is the introduction of Multi-agent Inverse Factorized Q-learning (MIFQ), a novel algorithm that employs mixing networks to aggregate decentralized Q functions for centralized learning and uses hyper-networks to generate weights for mixing networks [51]. This approach has demonstrated superior performance compared to baseline algorithms in various multi-agent environments, including SMACv2, Gold Miner, and Multi Particle Environments. MIFQ enables efficient and stable learning in cooperative multi-agent settings, and its objective function exhibits convexity within the Q function space under certain conditions.
@@
Another significant development in the field focuses on scaling laws for imitation learning in single-agent games. This research investigates the impact of scaling up model and data size on imitation learning performance, particularly in Atari games and NetHack [28]. By using Behavioral Cloning (BC) to imitate expert policies, the study reveals that imitation learning loss and mean return follow clear power law trends with respect to FLOPs. Importantly, loss and mean return are highly correlated, indicating that improvements in loss predictably translate to improved performance. The research demonstrates that scaling up model and data size can provide significant improvements in agent performance, with the scaled-up approach surpassing prior state-of-the-art by 1.5x in all settings for NetHack.
@@
In the realm of multi-agent systems, the Multi-Agent Adversarial Interaction Priors (MAAIP) approach adapts Multi-Agent Generative Adversarial Imitation Learning (MAGAIL) for modeling interactions between agents [20]. This method introduces new objectives for training the system and models self and opponent observations separately. MAAIP has proven effective for learning interactive behaviors between multiple agents and can be applied to scenarios where agents need to adapt to each other’s actions. This approach demonstrates potential for improving imitation learning in competitive or cooperative multi-agent settings.
@@
These recent advancements collectively highlight the importance of scaling, efficient centralized learning in decentralized execution settings, and modeling agent interactions for improved performance in complex environments. As the field of imitation learning continues to evolve, these contributions pave the way for more sophisticated and effective agent behaviors in both single-agent and multi-agent systems.
Imitation learning has found a wide range of applications in robotics, demonstrating its versatility and effectiveness in enabling robots to perform complex tasks. This section delves into three primary areas where imitation learning has been significantly applied: manipulation tasks, locomotion and navigation, and human-robot interaction.

### v2

# AI-Guided Molecular Simulations in VR: Exploring Strategies for Imitation Learning in Hyperdimensional Molecular Systems

@@ Abstract
Molecular dynamics (MD) simulations are a crucial computational tool for researchers to understand and engineer molecular structure and function in areas such as drug discovery, protein engineering, and material design. Despite their utility, MD simulations are expensive, owing to the high dimensionality of molecular systems. Interactive molecular dynamics in virtual reality (iMD-VR) has recently emerged as a ‘human-in-the-loop’ strategy for efficiently navigating hyper-dimensional molecular systems. By providing an immersive 3D environment that enables visualization and manipulation of real-time molecular simulations running on high-performance computing architectures, iMD-VR enables researchers to reach out and guide molecular conformational dynamics, in order to efficiently explore complex, high-dimensional molecular systems. Moreover, iMD-VR simulations generate rich datasets that capture human experts’ spatial insight regarding molecular structure and function. This paper explores the use of researcher-generated iMD-VR datasets to train AI agents via imitation learning (IL). IL enables agents to mimic complex behaviours from expert demonstrations, circumventing the need for explicit programming or intricate reward design. In this article, we review IL across robotics and Multi-agents systems domains which are comparable to iMD-VR, and discuss how iMD-VR recordings could be used to train IL models to interact with MD simulations. We then illustrate the applications of these ideas through a proof-of-principle study where iMD-VR data was used to train a CNN network on a simple molecular manipulation task – namely, threading a small molecule through a nanotube pore. Finally, we outline future research directions and potential challenges of using AI agents to augment human expertise in navigating vast molecular conformational spaces.
keywords:
molecular dynamics, virtual reality, imitation learning,
@@ 1 Introduction
Molecular dynamics (MD) simulations are a powerful tool for studying the structure, dynamics, and interactions of molecular systems. However, generating conformational ensembles and sampling rare events, (e.g., protein-ligand binding or protein conformational changes), remains challenging due to high computational costs and the complexity of the associated energy landscapes [1]. Interactive molecular dynamics in virtual reality (iMD-VR) has recently emerged as a promising approach to address these challenges by leveraging human intuition during real-time MD simulations within an immersive 3D environment [2]. In iMD-VR, users can directly manipulate and steer molecular systems using natural hand motions, applying forces to drive physically-relevant rare events such as conformational changes and ligand binding/unbinding [3]. This human-in-the-loop approach leverages the human’s innate ability for 3D spatial reasoning and manipulation, enabling a user to intuitively explore complex molecular landscapes. Recent studies have demonstrated the efficacy of iMD-VR in recreating crystallographic binding poses for protein-ligand systems [4, 5, 6] and generating important reactive pathways [7]. These interactive simulations capture valuable conformational data that can be challenging to obtain through conventional MD alone, thus offering new opportunities for applications such as training machine learning and investigating reaction mechanisms.
@@
Imitation learning (IL) is a machine learning paradigm where algorithms extract knowledge from expert demonstrations by either directly imitating the desired behaviour or learning the expert’s underlying reward function [8]. This approach allows machines to learn and replicate human actions, making it particularly valuable when manually designing reward functions is challenging, such as in noisy environments where it is easier to demonstrate desired behaviour than to engineer it. Learning from demonstration is often used interchangeably with IL in the literature [9, 8], though some work distinguishes between offline approaches using pre-collected demonstrations versus interactive IL, where learners can actively query experts for demonstrations during training. The literature systematizes IL approaches primarily based on the type of information available from expert demonstrations[10]. Learning from Demonstration (LfD) represents the traditional approach that involves both action and state supervisions, providing complete expert guidance. In contrast, Learning from Observations (LfO) is a more practical variant that only has access to state-only demonstrations, making it more challenging due to incomplete expert guidance but enabling the use of previously inapplicable resources such as videos. This approach has been particularly influential in the field of robotics, where it has been used to teach robots complex tasks without the need for explicit programming. By observing human demonstrations, robots can learn to perform a variety of actions, ranging from simple manipulations to complex, multi-step procedures [11, 12, 13, 14]. Learning from observation differs from other types of machine learning such as reinforcement learning, where an explicit reward function needs to be defined in advance or fine-tuned during training. IL offers the ability to learn a mapping of observations to actions performed by an expert in demonstrations, making it particularly well-suited to domains for which specifying a reward function is challenging or where human expertise can be leveraged. The versatility of IL has also sparked interest in its application beyond robotics [15], such as in MD [16], where it could potentially streamline the process of simulating and understanding complex molecular interactions.
@@
A policy is a mapping from observations or states of the environment to actions that the agent should take. IL often requires a large number of demonstrations to effectively learn it, especially for complex tasks. Collecting a sufficiently large and diverse dataset of human demonstrations can be challenging for various reasons [14]. One challenge is that human behaviour is often multi-modal—there are many valid ways to perform a task. Standard IL approaches may average out these modes and learn a sub-optimal policy. Capturing and replicating diverse human behaviours is an open challenge [17], mainly because robustly capturing all relevant aspects of human demonstrations can be difficult due to sensor limitations, occlusions, etc [18]. This is especially true when collecting data outside of lab settings ‘in the wild’. Therefore there is a growing need for more large-scale, open datasets of human demonstrations on standardized tasks in order to facilitate reproducible research and benchmark IL algorithms [17, 19].
@@
Virtual reality (VR) presents a novel and immersive platform for enhancing the capabilities of IL [14, 20], where VR is combined with high-performance computing to provide an interactive environment in which researchers can manipulate molecular structures in real-time. The intuitive and engaging nature of VR could transform how scientists interact with molecular simulations, making it easier to collect data, hypothesize, and test the dynamics of molecular systems [4, 21].
@@
This paper aims to review IL’s current applications in various domains, including key concepts like BC and GAIL. Particular focus is given to its potential in iMD-VR. We aim to provide a comprehensive review of the existing literature on IL, identify the benefits and challenges associated with its use, and propose innovative ways VR could serve as a data creation and collection platform for MD. By bridging the gap between IL and iMD-VR, we hope to open new avenues for research and application in the field of interactive MD, ultimately contributing to advancements in scientific understanding and technological development.
@@ 2.1 Molecular vizualisation
Virtual reality (VR) provides researchers with natural, intuitive 3D interfaces to view and interact with complex molecular structures in a way that is not facilitated with traditional 2D interfaces. This can enhance the researcher’s understanding of complex 3D molecular arrangements and interactions, which is essential for insight. Furthermore, VR can enhance scientific collaboration by providing shared virtual environments, which are accessible over the internet, thus enabling collaboration across physical distances.
There exist several programs for the visualization of molecular simulations in VR, e.g., UnityMol[22], and the commercial software Nanome[23]. These programs provides a collaborative virtual environment in which users can visualize and manipulate molecular structures in stereoscopic 3D. Researchers can analyse the spatial arrangement of molecules, measure distances between atoms, and dock ligands into protein binding pockets using natural hand gestures [24]. Other examples of software include ProteinVR [25], a web-based application that works across desktop, mobile, and VR platforms to visualize structural biology in 3D, and Molecular Rift [26], which provides controller-free manipulation of molecules using intuitive hand gestures.
By immersing users in 3D virtual environments, VR enables intuitive exploration of complex biomolecular systems that traditional 2D screens can not facilitate. The application of VR to molecular visualisation unlocks several key benefits. First, it provides researchers with natural, intuitive 3D interfaces to view and interact with complex molecular structures. Second, VR enables collaborative molecular modelling in shared virtual environments, enhancing scientific teamwork. Finally, the stereoscopic depth perception and wide field of view in VR leads to an enhanced spatial understanding of 3D molecular arrangements [27].
@@ 2.2 Interactive molecular simulations
Recent advancement in computational power and improving performance of graphical processing units has facilitated not only the visualization of molecules in VR, but also real-time interactivity. One example is NanoVer [28, 3, 29], an open-source program developed by Glowacki et al. for performing iMD-VR. Using VR controllers, users can apply forces directly to MD simulations in real-time to drive important chemical events such as ligand binding and conformational changes.
Several aspects of NanoVer make it a suitable software for use in an iMD-VR+IL workflow. Firstly, NanoVer delivers quantitative chemical information on-the-fly such as potential energy, kinetic energy, and work done on the system during user interactions. These quantities provide physically motivated metrics with which to train IL models. Narupa [3, 29] (the precursor to NanoVer) has already been used to investigate reaction pathways [7], ligand binding poses [4, 5], and thermodynamic properties such as binding free energies [6], which validates the use of this iMD-VR framework as a research-grade tool. To our best knowledge, this is the only open-source iMD-VR software that has proven utility in a chemistry research context. Furthermore, the program offers comprehensive documentation, active ongoing development, and compatibility with the latest commercial VR headsets.
@@
Figure 1: Showing an iMD-VR user docking and undocking the drug oseltamivir from the H7N9 neuraminidase protein (see text for further detail) [3]
@@
An example of this is demonstrated in Figure 1. The top panel illustrates a researcher using iMD-VR to explore the binding and unbinding pathways of the drug oseltamivir, also known as tamiflu, (shown in magenta) with the H7N9 neuraminidase protein. The bottom panel shows snapshots of the molecular system at three time-points in the simulation: (1) oseltamivir bound to the active site of neuraminidase, (2) the molecular system after the researcher has undocked oseltamivir from the binding pocket of neuraminidase, and (3) the protein-ligand complex, where oseltamivir has been re-docked by the researcher after interactively exploring potential binding modes.
2.3 Molecular data structures in iMD-VR
@@
In essence, an MD simulation is a time series consisting of a set of frames of the atomic positions of a molecular system, which can be viewed frame-by-frame as a ‘trajectory’ (an example is shown in Table 1). In iMD-VR, these trajectories are generated and visualized on the fly. Each frame contains information about the system (e.g. the temperature and energy), and the atoms contained within it (e.g. their position, velocities and element type). The iMD-VR program NanoVer uses a key-value system to store and communicate the data associated with these frames.
@@
Table 1: Dataframe of the first frame from a recording for the nanotube task
atom name frame index coordinates user forces
@@
[…]
NanoVer streams data in real-time in the form of two dictionaries: (a) the frames, containing information about the simulation, such as the atomic positions, velocities, energies, and forces; and (b) the ‘shared state’, consisting of synchronized information about avatar positions and user interactions such as controller positions and button pushes. The NanoVer server uses these streams to manage and process how users interact with MD simulation in VR. Moreover, it can record these data, thus enabling post-hoc analysis and playback of sessions. These recordings can be loaded onto the server, which then sends the recorded streams (synchronized using timestamps) to the clients as if they were real-time simulation streams. In this case, NanoVer acts purely as a molecular visualizer, affording the user control over the position/rotation/scale of the simulation and providing typical playback features (play/pause/restart). During playback, the user cannot apply forces to the molecular system. NanoVer recordings can be imported and analysed using a Python script with the MDanalysis package[30, 31].
NanoVer offers developers and users the possibility of plugging in a selection of MD simulation software. One of the prototypical examples used for NanoVer demonstrations is the simulation of a methane molecule and a carbon nanotube, a molecular system that memics biomolecular channels that act as molecule-selective filters[32].
Figure 2: Trajectory depicting threading of a Methane molecule through a carbon Nanotube
In this simulation, users can thread a methane through a nanotube (Figure 2). The molecular system comprises 65 atoms: 60 carbons for the nanotube (labeled C1–C60), and 1 carbon and 4 hydrogens for methane (labeled C61 and H1–H4). Table 1 shows an example of the data collected for this system. Such systems, where users try to reach a specific goal (e.g. achieve a specific conformation, follow specific reaction coordinates, etc.), will be referred to as “molecular tasks”.
The data frame shown in Table 1 has 4 columns: atom name, frame index, coordinates, and user forces. Here, atom name is the atom’s label, frame index is the simulation frame index, coordinates contain the (x,y,z) positions in nanometres, and forces contain the (x,y,z) components of the forces in kilojoules per mole per nanometre (Fx, Fy, Fz) applied by the user on the specified atom.
@@
A more sophisticated molecular task is the one shown in Figure 3. This task involves tying a knot in a 17-alanine molecule. The user is presented with the molecule in native linear form and asked to form a knot by manipulating both ends. The systems illustrated in Figs 2 and 3 are demonstration systems involving only small molecules. The software can be used on larger, more complex molecular systems such as protein systems.
@@ 3.1 Recent works in literature
The recent surge of research on IL [14, 33, 9, 10] where policies are acquired by mimicking expert demonstrations instead of optimizing explicit reward definitions has generated a rich methodological landscape and a broad spectrum of applications. To orient this work, the literature review is structured into two subsections. The first contrasts the use of IL in single-agent and multi-agent systems, which studies on one side a solitary learner acting in an otherwise stationary environment, and on the other side multi-agent IL, where several learners must coordinate and reason over multiple interacting components; recent advances highlight how complex behaviours can emerge from simpler local interactions. This is particularly relevant for IL in iMD-VR to understanding how coordinated molecular manipulations can lead to desired conformational changes or structural transitions. The second subsections examine why IL has become a cornerstone of modern robotics: surveys and empirical studies show how it delivers sample-efficient, safety-aware skill acquisition for tasks ranging from autonomous driving to dexterous manipulation. The spatial reasoning capabilities developed in robotics are particularly relevant to molecular systems to developed imitation learning approaches for complex manipulation tasks that require understanding three-dimensional spatial molecular relationships. Together, these perspectives lay the conceptual and practical groundwork for leveraging insights from both algorithmic strands to inform the application of IL within iMD-VR. Both robotics and MAS address challenges that are central to iMD-VR applications: handling high-dimensional state spaces, learning from expert demonstrations, coordinating multiple interacting components, and achieving robust performance under uncertainty. The high-dimensional sampling problem that makes molecular dynamics computationally expensive parallels the high-dimensional control challenges in both robotics manipulation and multi-agent coordination. The human-in-the-loop aspect of iMD-VR also benefits from both domains - robotics provides frameworks for learning from human demonstrations and teleoperation, while MAS offers insights into human-agent interaction and coordination strategies.
By reviewing both robotics and MAS applications, we can leverage the complementary strengths of both fields: robotics’ expertise in manipulation and spatial reasoning, and MAS’s expertise in coordination and distributed control, creating a more comprehensive foundation for advancing molecular manipulation capabilities in virtual reality environments.
Whereas in single-agent systems, IL algorithms operate by training an individual agent to mimic expert demonstrations, focusing on state-action mapping without considering interactions with other learning entities, in multi-agent systems, imitation learning becomes significantly more complex as algorithms must simultaneously learn individual agent policies and coordination mechanisms from demonstrations of multiple interacting experts. Seeing that NanoVer[28] supports multi-user interactions and enables multiple persons to be in the same VR molecular environment, this multi-user capability directly aligns with multi-agent IL principles, as users can:
  • Generate Coordinated Demonstrations: Expert molecular manipulation teams can provide demonstrations of collaborative molecular dynamics tasks, showing how different users coordinate to achieve complex molecular manipulations.
  • Learn Coordinated Policies: Multi-agent IL algorithms can learn from these multi-user demonstrations to understand both individual manipulation strategies and the coordination patterns that emerge during collaborative molecular exploration.
@@
In recent years, IL in agents and multi-agent systems has seen significant advancements. One notable contribution is the introduction of Multi-agent Inverse Factorized Q-learning (MIFQ) [34], a novel algorithm that employs mixing networks to aggregate decentralized Q functions for centralized learning and uses hyper-networks to generate weights for mixing networks [35]. This approach has demonstrated superior performance compared to baseline algorithms in various multi-agent environments, including SMACv2 [36], Gold Miner [37], and Multi Particle Environments. MIFQ enables efficient and stable learning in cooperative multi-agent settings.
@@
Another significant development in the field focuses on scaling laws for IL in single-agent games. Paine et al. [38] investigate the impact of scaling up model and data size on IL performance, particularly in Atari games and NetHack [39]. By using Behavioural Cloning (BC) to imitate expert policies, they reveal that IL loss and mean return follow clear power law trends with respect to FLOPs. Importantly, loss and mean return are highly correlated, indicating that improvements in loss predictably translate to improved performance. Their research demonstrates that scaling up model and data size can provide significant improvements in agent performance, with the scaled-up approach surpassing prior state-of-the-art by 1.5x for NetHack.
@@
In the realm of multi-agent systems, the Multi-Agent Adversarial Interaction Priors (MAAIP) [40] approach adapts Multi-Agent Generative Adversarial IL (MAGAIL) for modelling interactions between agents. Younes et al. [40] introduce new objectives for training the system and modelling self and opponent observations separately. MAAIP has proven effective for learning interactive behaviours between multiple agents and can be applied to scenarios where agents need to adapt to each other’s actions. This approach demonstrates potential for improving IL in competitive or cooperative multi-agent settings.
@@
These recent advancements collectively highlight the importance of scaling, efficient centralized learning in decentralized execution settings, and modelling agent interactions for improved performance in complex environments. As the field of IL continues to evolve, these contributions pave the way for more sophisticated and effective agent behaviours in both single-agent and multi-agent systems.